# Humidity-Aware Model Predictive Control for Residential Air Conditioning: A Field Study

### v1

@@ 5.1 Controller behavior
The different control system implementations were tested continuously from July 10th to September 7th 2023. Of these 60 days, seven were lost due to intermittent thermostat Wi-Fi issues, seven were lost due to controller issues (badly tuned comfort price, communication issues), and eight were reverted back to non-MPC control. Of the remaining 38 days, 24 used cost-reducing control, while 14 performed power-limiting control. For the power-limiting MPC, the re-calibrated thermal circuit model in Section 3.4 was used, while for the cost-reducing MPC, the original values presented in Section 3.3 were used. Across the 38 testing days, the “sensible” and “latent” controller formulations were tested in an one-day-on, one-day-off fashion, with a total of 21 “latent” days and 17 “sensible.”Fig. 10 shows the performance of the controller on a typical cost-reducing MPC day. The set-point typically increases during the hot afternoon, when the COP is low, to save electricity. The set-point typically returns to the occupants’ preference overnight and through the morning.
@@ 5.2 Thermal comfort
The supervisory control system’s thermal comfort performance was assessed in two ways. First, a web form was offered for occupants to record any thermal discomfort they felt. During the total duration of testing, only one response was received, and that was of mild dissatisfaction (as per [26]). Second, a thermal comfort software package [27] was used to generate a PPD time series based on indoor temperature measurements and other inputs to the PPD model, such as occupants’ estimated clothing and activity levels. This process yielded a time-average PPD of 7.2% over the MPC test hours and 8.5% over the test hours without MPC. Practitioners typically consider $\leq 10$% PPD acceptable [26]. The PPD was in the slightly uncomfortable region for a total duration of only two hours across the MPC testing days.
@@ 5.4 Power limiting
“Sensible” 7 2.96 54 0.8
“Latent” 7 3.01 11 (80% reduction) 0.3 (63% reduction)
@@
The results of testing over a hot week are shown in Fig. 14. The “latent” formulation predicted the true electrical peaks more accurately, allowing it to drift the required amount to minimize the power use over the window. Fig. 15 shows the performance of the MPC algorithms in mitigating power peaks. MPC with the “latent” formulation reduces both the the magnitude and frequency of constraint violation. “Sensible” MPC violated the power limit between 4 and 8 PM for an average of 54 minutes per day, with a conditional-mean magnitude of 0.8 kW. By contrast, “latent” MPC violated the constraint for an average of 11 minutes per day, with a conditional-mean magnitude of 0.3 kW. Relative to “sensible” MPC, “latent” MPC reduced the daily mean duration of constraint violation by 80% (from 54 to 11 minutes) and the conditional-mean magnitude of constraint violation by 63% (from 0.8 to 0.3 kW).
@@ 5.5 Comparison of humidity modeling approaches
Table 1 compares the performance of MPC with the “sensible” and “latent” humidity modeling approaches, for both energy efficiency and power limiting. In this table, the weather-normalized energy is defined as the ratio of the cumulative daily heat pump electrical power to the cumulative daily indoor-outdoor temperature difference, averaged over $N$ total testing days:
@@
Table 1 shows that “sensible” MPC and “latent” MPC performed similarly for energy efficiency, with weather-normalized energies of 2.32 and 2.34 kWh/C, respectively, for cost-reducing MPC. This finding corroborates previous observations [47, 49] that model accuracy has relatively little impact on energy savings. However, “latent” MPC significantly outperformed “sensible” MPC for power-limiting control, reducing the mean duration of constraint violation by 80% and the conditional-mean violation magnitude by 63%. These results suggest that model accuracy matters more when objectives or constraints depend sensitively on the magnitude of power use.
@@ 6.1 Practical considerations
Impact of assumptions. One assumption underlying this work was the numerical value of the SHR used for the “sensible” MPC formulation. After testing, it was found that the conditions were more humid than expected: the SHR actually averaged 0.79, while the “sensible” MPC assumed an SHR of 0.86. It is challenging to select an appropriate SHR value, as (1) it can change from house to house based on floor space and envelope characteristics, (2) it is highly dependent on the weather and on internal moisture generation driven by occupant behavior, and (3) it can vary based on the capacity of the cooling equipment and heat exchanger characteristics. Typically, to avoid these issues, researchers do not model the latent load or the indoor humidity in experiments of supervisory control [7, 8, 49]. Another limitation of this study lies in the testing of the controller in a single climatic zone and a single household. More data are needed to assess performance under varying occupant preferences and housing types across different climates.
@@ 7 Conclusion
This paper presented the first supervisory control experiment for residential air conditioning that accounted for time-varying humidity effects. It developed novel data-driven time-series forecasting methods to predict indoor humidity conditions. This paper tested MPC with two humidity models: a “sensible” model that assumed a constant sensible heat ratio, and a “latent” model that predicted a time-varying sensible heat ratio based on dynamic predictions of the return air wet-bulb temperature. Each modeling approach was tested in two MPC formulations: one that sought to minimize energy costs, and one that sought to constrain electrical power below a utility-specified limit during a demand response event. “Sensible” and “latent” MPC performed similarly for energy cost reduction, but “latent” MPC performed much better for power limiting. These results corroborate a finding of other researchers – that energy cost savings depend weakly on model accuracy – but suggest that model accuracy is more important when minimizing or constraining peak power demand.

### v2

@@ 5.1 Controller behavior
The different control system implementations were tested continuously from July 10th to September 7th 2023. Of these 60 days, seven were lost due to intermittent thermostat Wi-Fi issues, seven were lost due to controller issues (badly tuned comfort price, communication issues), and eight were reverted back to non-MPC control. Of the remaining 38 days, 24 used cost-reducing control, while 14 performed power-limiting control. For the power-limiting MPC, the re-calibrated thermal circuit model in Section 3.4 was used, while for the cost-reducing MPC, the original values presented in Section 3.3 were used. Across the 38 testing days, the sensible and latent controller formulations were tested in an one-day-on, one-day-off fashion, with a total of 21 latent days and 17 sensible. Fig. 11 shows the performance of the controller on a typical cost-reducing MPC day. The set-point typically increases during the hot afternoon, when the COP is low, to save electricity. The set-point typically returns to the occupants’ preference overnight and through the morning.
@@ 5.2 Thermal comfort
The supervisory control system’s thermal comfort performance was assessed in two ways. First, a web form was offered for occupants to record any thermal discomfort they felt. During the total duration of testing, only one response was received, and that was of mild dissatisfaction (as per [29]). Second, a thermal comfort software package [30] was used to generate a PPD time series based on indoor temperature measurements and other inputs to the PPD model, such as occupants’ estimated clothing and activity levels. This process yielded a time-average PPD of 7.2% over the MPC test hours and 8.5% over the test hours without MPC. Practitioners typically consider $\leq 10$% PPD acceptable [29]. The PPD was in the slightly uncomfortable region for a total of two hours across the MPC testing days. While these PPD results are encouraging, they are not definitive. PPD may not accurately reflect comfort under highly transient conditions [54]. Furthermore, the PPD calculations here assumed constant air flows, activity levels, and clothing levels; in reality, these varied somewhat over the experiments. Nevertheless, the combined results of the PPD calculations and occupant questionnaires suggest that MPC maintained satisfactory comfort with both the latent and sensible MPC models.
@@ 5.4 Power limiting
Sensible 7 2.96 54 0.8
Latent 7 3.01 11 (80% reduction) 0.3 (63% reduction)
@@
The results of testing over a hot week are shown in Fig. 15. The latent formulation predicted the true electrical peaks more accurately, allowing it to drift the required amount to minimize the power use over the window. Fig. 16 shows the performance of the MPC algorithms in mitigating power peaks. MPC with the latent formulation reduces both the the magnitude and frequency of constraint violation. sensible MPC violated the power limit between 4 and 8 PM for an average of 54 minutes per day, with a conditional-mean magnitude of 0.8 kW. By contrast, latent MPC violated the constraint for an average of 11 minutes per day, with a conditional-mean magnitude of 0.3 kW. Relative to sensible MPC, latent MPC reduced the daily mean duration of constraint violation by 80% (from 54 to 11 minutes) and the conditional-mean magnitude of constraint violation by 63% (from 0.8 to 0.3 kW).
@@ 5.5 Comparison of humidity modeling approaches
Table 1 compares the performance of MPC with the sensible and latent humidity modeling approaches, for both energy efficiency and power limiting. In this table, the weather-normalized energy is defined as the ratio of the cumulative daily heat pump electrical power to the cumulative daily indoor-outdoor temperature difference, averaged over $N$ total testing days:
@@
Table 1 shows that sensible and latent MPC performed similarly for energy efficiency, with weather-normalized energies of 2.32 and 2.34 kWh/C, respectively, for cost-reducing MPC. However, latent MPC significantly outperformed sensible MPC for power-limiting control, reducing the mean duration of constraint violation by 80% and the conditional-mean violation magnitude by 63%. These results suggest that model accuracy matters more when objectives or constraints depend sensitively on the predicted magnitude of power use.
These results, alongside prior literature [10, 15, 56], suggest that accurate humidity modeling can improve performance from some residential HVAC control tasks. For other tasks, however, good performance can be achieved without modeling humidity accurately, or indeed modeling it at all [8, 57]. For example, Table 1 shows that despite its oversimplified humidity model, sensible MPC still reduced the mean duration of violating the power-limiting constraint by 40% relative to the non-MPC baseline, from 90 minutes of violation to 54.
@@ 6.1 Practical considerations
Impact of assumptions. One assumption underlying this work was the numerical value of the SHR used for the sensible MPC formulation. After testing, it was found that the conditions were more humid than expected: the SHR actually averaged 0.79, while the sensible MPC assumed an SHR of 0.86. It is challenging to select an appropriate SHR value, as (1) it can change from house to house based on floor space and envelope characteristics, (2) it is highly dependent on the weather and on internal moisture generation driven by occupant behavior, and (3) it can vary based on the capacity of the cooling equipment and heat exchanger characteristics. Typically, to avoid these issues, researchers do not model the latent load or the indoor humidity in experiments of supervisory control [7, 8, 28]. Another limitation of this study lies in the testing of the controller in a single climatic zone and a single household. More data are needed to assess performance under varying occupant preferences and housing types across different climates.
@@ 7 Conclusion
This paper presented the first supervisory control experiment for residential air conditioning that accounted for time-varying humidity effects. It developed novel data-driven time-series forecasting methods to predict indoor humidity conditions. This paper tested MPC with two humidity models: a sensible model that assumed a constant SHR, and a latent model that predicted a time-varying SHR based on dynamic predictions of the return air wet-bulb temperature. Each modeling approach was tested in two MPC formulations: one that sought to minimize energy costs, and one that sought to constrain electrical power below a utility-specified limit during a demand response event. sensible and latent MPC performed similarly for energy cost reduction, but latent MPC performed much better for power limiting. These results corroborate a finding of other researchers – that energy cost savings depend weakly on model accuracy – but suggest that model accuracy is more important when minimizing or constraining peak power demand.